OpenStudy (anonymous):

Which of these are the same as 64:512?? A. (8/4)^2 B. (4/8)^3 C. (8/4)^3 D. (4/8)^3

hartnn (hartnn):

64=4^3 and 512=8^3 does this help?

OpenStudy (anonymous):

So it would be B?

hartnn (hartnn):

yup:)

hartnn (hartnn):

but isn't that same as D???

OpenStudy (anonymous):

D is supposed to be ^2 sorry

hartnn (hartnn):

then it would be B

OpenStudy (anonymous):

Thanks! (: